A cast iron plate is described below:
Circular.
Diameter: 4' 8"
Thickness; 1"
Has 26 1.5" holes, evenly spaced.
Determine weight.
Cast iron weighs .26 lbs. / in.^3

Plate is a cylinder.

Volume = pi * r^2 * h
3.1416 * 24.333 * 24.333 * 1 = 1860.1253 in.^3 (total vol.)
3.1416 * .75 * .75 * 1 * 26 = 49.94 = 50 in.^3
1860 - 50 = 1810 (net vol.)

1810 * .26 = 470.6 = 471 lbs.
Answer is larger. Where am I right?


Answer by Alan3354(69443)

A cast iron plate is described below:
Circular.
Diameter: 4' 8"
Thickness; 1"
Has 26 1.5" holes, evenly spaced.
Determine weight.
Cast iron weighs .26 lbs. / in.^3

Plate is a cylinder.

Volume = pi * r^2 * h
3.1416 * 24.333 * 24.333 * 1 = 1860.1253 in.^3 (total vol.)
3.1416 * .75 * .75 * 1 * 26 = 49.94 = 50 in.^3
1860 - 50 = 1810 (net vol.)

1810 * .26 = 470.6 = 471 lbs.
Answer is larger. Where am I right?
-----
You're right that it's a cylinder.
==================
Volume = pi*r^2*h
r = 28 inches, not 24 1/3